OpenStudy (anonymous):

Solve the following equation 3^x + 1 - 3^x = 54 how do you solve if you can't make 54 into a base of 3????

OpenStudy (anonymous):

since 54/3 = 18; that might make it work

OpenStudy (anonymous):

let me work it out

OpenStudy (asnaseer):

is this your question?\[3^x+1-3^x=54\]

OpenStudy (asnaseer):

or is it:\[3^x+(1-3)^x=54\]or something else?

OpenStudy (anonymous):

oh.. i should make it clearer 3^(x + 1) - 3^x = 54

OpenStudy (asnaseer):

ok, so we have:\[3^{x+1}-3^x=54\]therefore:\[3*3^x-3^x=54\]therefore:\[2*3^x=54\]can you solve from here?

OpenStudy (anonymous):

Yes!! thank you that makes a lot of sense!!

OpenStudy (asnaseer):

yw